OpenStudy (anonymous):

derivative of 1/bx? I know.... so easy, just bear with me @Calculus1

OpenStudy (anonymous):

1/bx=bx^-1 y'= -1(bx)^-2 * (b+x)

OpenStudy (eyust707):

no not b+x

OpenStudy (anonymous):

I know that isn't right ^^^ ... why in the world am I messing up! It's so simple too... Maybe I need to take a break. Been studying for 6 hours straight lol

OpenStudy (anonymous):

chain rule?

OpenStudy (eyust707):

yep

OpenStudy (eyust707):

what the derivitive of bx?

OpenStudy (anonymous):

deriv of first times second + deriv of second times first 1x+1b

OpenStudy (eyust707):

chain rule says "take the derivative of the outside then multiply it by the derivative of whats on the inside"

OpenStudy (eyust707):

that is the product rule you are using... typically b is assumed to be a constant

OpenStudy (anonymous):

\[\frac{d}{dx}\frac{1}{bx}=\frac{1}{b}\times \frac{d}{dx}\frac{1}{x}=\frac{1}{b}\times \frac{-1}{x^2}=-\frac{1}{bx^2}\]

OpenStudy (anonymous):

^^k. Brain fart. TY

OpenStudy (anonymous):

constant is not important. ignore it and multiply at the end